OpenStudy (anonymous):

The U.S. Weather Service reports that in a certain northern city it rains 40 days and snows 50 days in the winter. However, it rains and snows on only 10 of those days. Based on this information, what is the probability that it will rain or snow in that city on a particular winter day? (Assume that there are 90 days of winter.)

OpenStudy (anonymous):

probability of rain or snow=(40/90)+(50/90)-(10/90)=8/9